the initial velocity of a particle moving along a straight line is 10m/s and its retardation is 2m/s^2. The distance moved by the particle in the 5th second will be

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
initial velocity=u=10m/s
Acceleration = -Retardation= -2 m/s2
Velocity after 4 seconds sice we have to find the distance travelled in 5th second= V5=u+at
=>V5= 10-2x4 =8m/s

Distance moved by the particle in the fifth second = Displacement of the particle in the fifth second (since they are moving in straight line)
=>Displacement=ut+at2/2
                           =2x1-2x1/2
                           =1m
Distance moved=1m.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)